Pentane is burnt with 96000 cm3 of oxygen gas at room temperature and pressure.


a) write the balance chemical equation for the reaction.


b) calculate the volume of carbon dioxide in cubic centimeters produce in the reaction at room temperature and pressure


c) calculate the mass of zinc needed to react with 50 cubic centimeters of 0.8 M of Hydrochloric acid

1
Expert's answer
2022-05-05T12:00:03-0400

a) C5H12 + 8O2 = 5CO2 + 6H2O

b) Molar mass of Oxygen = 31.998

96000/31.998 = 3000.19m

Molar mass of CO2 = 44.01 g/mol

V(CO2) = 3000.19× 44.01× 5/8 = 82523.98 cm3

c) Zn + 2HCl = ZnCl2 + H2

n(HCl) = 0.8 m • 0.05 L = 0.04 moles

m(Zn) = (m(HCl)•Mr(Zn)) / 2Mr(HCl)

m(Zn) = (1.46 g • 65g/mol) / 2•36.5 g/mol = 1.3 g